# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

### v1

@@ Abstract
Kolmogorov-Arnold networks (KANs) are a remarkable innovation consisting of learnable activation functions with the potential to capture more complex relationships from data. Although KANs are useful in finding symbolic representations and continual learning of one-dimensional functions, their effectiveness in diverse machine learning (ML) tasks, such as vision, remains questionable. Presently, KANs are deployed by replacing multilayer perceptrons (MLPs) in deep network architectures, including advanced architectures such as vision Transformers (ViTs). In this paper, we are the first to design a general learnable Kolmogorov-Arnold Attention (KArAt) for vanilla ViTs that can operate on any choice of basis. However, the computing and memory costs of training them motivated us to propose a more modular version, and we designed particular learnable attention, called Fourier-KArAt. Fourier-KArAt and its variants either outperform their ViT counterparts or show comparable performance on CIFAR-10, CIFAR-100, and ImageNet-1K datasets. We dissect these architectures’ performance and generalization capacity by analyzing their loss landscapes, weight distributions, optimizer path, attention visualization, and spectral behavior, and contrast them with vanilla ViTs. The goal of this paper is not to produce parameter- and compute-efficient attention, but to encourage the community to explore KANs in conjunction with more advanced architectures that require a careful understanding of learnable activations. Our open-source code and implementation details are available on: subhajitmaity.me/KArAt
Multi-Head Self-Attention, Vision Transformers, Kolmogorov-Arnold Network, Kolmogorov-Arnold Transformers, Kolmogorov-Arnold Attention
Figure 1: Model parameters vs. Top-1 Test accuracy in ImegeNet-1K training of vanilla ViTs (Dosovitskiy et al., 2020), Vision KAN (DeiT+KAN) by (Chen et al., 2024), ViT+KAN and Kolmogorov-Arnold Transformer (KAT) by (Yang & Wang, 2025).
@@ 1 Introduction
Artificial general intelligence has become a rapidly growing research direction, and Kolmogorov-Arnold Network (KAN) (Liu et al., 2024) marks a remarkable innovation to that. KANs with learnable activation functions can potentially capture more complex relationships and facilitate meaningful interaction between the model and human intuition. After training a KAN on a specific problem, researchers can extract the learned univariate functions the model uses to approximate complex multivariable functions. By studying these learned functions, researchers can gain insights into the underlying relationships from the data and refine the model. KANs exhibit state-of-the-art performance in finding symbolic function representations (Yu et al., 2024), continual learning of one-dimensional functions (Liu et al., 2025).
@@
KANs were integrated with neural network architectures, the primary of them being conventional MLPs or convolution neural networks (CNNs). E.g., (Ferdaus et al., 2024; Bodner et al., 2024; Drokin, 2024; Abueidda et al., 2025; Wang et al., 2025) combine KAN with CNNs, (Li et al., 2024) combine KAN with U-Net, etc. Interestingly, (Yu et al., 2024) claimed to make the first fairer comparison between KANs and MLPs on multiple ML tasks on small-scale datasets. By setting the same parameter counts in both networks, (Yu et al., 2024) concluded that KANs with B-Spline basis significantly underperform compared to their MLP counterparts in all ML tasks except symbolic formula representation; a similar claim reinstated in (Azam & Akhtar, 2024) for vision. But KAN’s exploration of more advanced architectures such as Transformers, remains limited. (Chen et al., 2024) replace the MLP layers inside the encoder blocks of data-efficient image Transformers (DeiTs) (Touvron et al., 2021) with KANs and proposed DeiT+KAN, (Yang & Wang, 2025) proposed two variants; ViT+KAN that replaces the MLP layers inside ViT’s encoder blocks, and Kolmogorov-Arnold Transformer (KAT), albeit similar to ViT+KAN but with a refined group-KAN strategy.
@@
While KANs were claimed to be not superior to simple MLPs in vision tasks (Yu et al., 2024; Azam & Akhtar, 2024), we were curious to find out how ViTs can see images through the lens of KANs. In that effort, Figure 1 demonstrates the performance of vanilla ViTs (ViT tiny, ViT small (Wu et al., 2022; Steiner et al., 2022), and ViT base (Dosovitskiy et al., 2020)) and their KAN counterparts on ImageNet-1K (Deng et al., 2009) training. While different variants of DeiT+KAN (Chen et al., 2024) and ViT+KAN (Yang & Wang, 2025) show a 5%–18% drop in the Top-1 accuracy compared to vanilla ViTs and DeiTs in ImageNet-1K, the KATs (Yang & Wang, 2025) show approximately 1%–2.5% gain in Top-1 accuracy compared to vanilla ViTs and DeiTs by keeping about the same number of parameters. These results show KANs require further investigation in more advanced architectures such as Transformers. Nevertheless, first, we want to understand why there is a discrepancy in these two seemingly equivalent implementations.
@@
In search for an answer, we realized that while DeiT+KAN and ViT+KAN replace the MLP layer with KAN in the Transformer’s encoder block, KAT implements a sophisticated group-KAN strategy that reuses a learnable function among a group of units in the same layer and chooses different bases for different encoder blocks. This strategy helps improve the accuracy by keeping almost similar parameter counts. Therefore, simply replacing MLPs with KANs might not guarantee better generalizability, but a properly designed KAN could. This perspective on the KANs’ performance in conjunction with Transformers opens many research questions. Importantly, we also note that the community remains put from designing a learnable multihead attention module, the heart of the Transformers. Therefore, in the rise of second-generation Transformers, such as Google’s TITAN (Behrouz et al., 2024) and SAKANA AI’s Transformer2 (Sun et al., 2025) that mimic the human brain, we want to ask: Is it worth deploying learnable multi-head self-attention to the (vision) Transformers?
Figure 2: (a) The standard softmax self-attention for $i^{\rm th}$ head in the $j^{\rm th}$ encoder block. (b) The Kolmogorov-Arnold Attention (KArAt) replaces the softmax with a learnable function $\Phi^{i,j}$. (c) The ideal KArAt is with an operator matrix, $\Phi^{i,j}$ with $N\times N$ learnable units that act on each row of ${\cal A}^{i,j}$. (d) While $\Phi^{i,j}\in\mathbb{R}^{N\times N}$ is impossible to implement due to computational constraints, our architecture uses an operator $\widehat{\Phi}^{i,j}$ with $N\times r$ learnable units $r\ll N$, followed by a linear projector with learnable weights $W\in\mathbb{R}^{r\times N}$.
@@
Finally, we note that there is a line of research that proposes efficient attention mechanisms in terms of sparse attention (Rahimian et al., 2024; Yun et al., 2020; Shi et al., 2021; Kovaleva et al., 2019; Zhang et al., 2021; Zaheer et al., 2020; Guo et al., 2019) or linear/kernelized attention (Han et al., 2025; Nguyen et al., 2023, 2021, 2022b; Lu et al., 2021; Nguyen et al., 2022a), to remedy computational and memory complexities of attention calculation. In contrast, this paper tries to understand how learnable multi-head self-attention (MHSA) modules can perform over regular self-attention used in ViTs, a defining technology in vision in the last six years. We list our contributions as follows:
@@
Designing a Learnable Attention Module for ViTs. Since language is significantly more complicated, less structured, and has a higher information entropy than an image, we limit this pilot study to the ViTs (Dosovitskiy et al., 2020) for image classification. In §3, we propose a general learnable Kolmogorov-Arnold multi-head self-attention mechanism or KArAt for vanilla ViTs that can operate on any basis. However, the computing and memory costs of training them motivated us to propose a more modular version, and we designed a particular MHSA mechanism using Fourier basis, called Fourier KArAt and its variants; see §4 for details.
@@
Benchmarking, Evaluation, and Analysis. In §5, we benchmark different variants of Fourier KArAt on CIFAR-10, CIFAR-100, (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009) datasets and evaluate their performance against vanilla ViTs—ViT tiny, ViT small (Wu et al., 2022; Steiner et al., 2022), and ViT base (Dosovitskiy et al., 2020). We dissect these architectures’ performance and generalization capacity by analyzing the loss landscapes, weight distributions, optimizer path, attention visualization, and spectral behavior and contrast them with vanilla ViTs.
@@ 2 Background
Multi-layer Perceptrons (MLPs) consist of $L$ layers, where the $l^{\rm th}$-layer is defined for input $x^{[l-1]}\in\mathbb{R}^{d_{l-1}}$ as
@@
By convention, $a^{[0]}=x^{[0]}$ denotes the input data. In MLP, the nonlinear activation functions, $\Phi(\cdot)$ are fixed (Bengio et al., 2017). For supervised tasks, given a training dataset $D$ with $N$ elements of the form (input, ground-truth) pairs, $\{(x_{i},y^{\star}_{i})\}_{i=1}^{N}$, the loss, $\mathcal{L}(X,Y^{\star}|\mathcal{W})=d_{Y}(a^{[L]},Y^{\star})$ (metric induced by the space $\mathbb{R}^{d_{L}}$) is calculated in the forward pass. The layerwise weights, $\{(W_{l},b_{l})\}_{l\in[L]}$ are learned by minimizing $\mathcal{L}(X,Y^{\star}|\mathcal{W})$.
Kolmogorov-Arnold Network (KAN) (Liu et al., 2025) is a neural network involving learnable activations parametrized by a chosen set of basis functions defined on a set of grid points or knots. The idea for this network stems from the Kolmogorov-Arnold Representation Theorem.
@@
(Kolmogorov, 1956) For any multivariate continuous function, $f:[0,1]^{n}\to\mathbb{R}$, there exists a finite composition of continuous single-variable functions, $\phi_{q,p}:[0,1]\to\mathbb{R},\Phi_{q}:\mathbb{R}\to\mathbb{R}$ such that
$f(x)=f(x_{1},x_{2},\dots,x_{n})=\sum\limits_{q=1}^{2n+1}\Phi_{q}\bigg{(}\sum\limits_{p=1}^{n}\phi_{q,p}(x_{p})\bigg{)}.$
Theorem 1 describes an exact finite expression using 2 layers, but, this can be extended to a multi-layer KAN; it is analogous to the Universal Approximation Theorem (Hornik et al., 1989).
@@
For input $x^{[0]}$, an $L$-layer KAN is given as $\text{KAN}(x^{[0]})=\Phi^{L}\circ\cdots\circ\Phi^{l}\circ\cdots\circ\Phi^{1}(x^{[0]})$. In each KAN layer, the activation function $\Phi$ is learnable, and ${\Phi^{[l]}=[\Phi_{ij}^{[l]}]}$ operates on ${x^{[l-1]}\in\mathbb{R}^{d_{l-1}}}$ to produce ${x^{[l]}\in\mathbb{R}^{d_{l}}},$ such that:
${x^{[l]}_{i}=\Phi^{[l]}_{i:}(x^{[l-1]})=\sum\limits_{j=1}^{d_{l-1}}\Phi^{[l]}_{ij}(x^{[l-1]}_{j})}.$
@@
Originally, (Liu et al., 2025) chose the B-spline basis functions. That is, the activation functions were defined to be $\phi(x)=w(b(x)+\text{spline}(x))$, where $b(x)=\text{silu}(x)=\frac{x}{1+e^{-x}}$ and $\text{spline}(x)=\sum\limits_{i}c_{i}B_{i}(x)$, with $B_{i}(\cdot)$ being one of the $k$-th degree B-spline basis functions parametrized by the $G$ grid points on a uniform grid, $[-I,I]$. In this case, the representation for each function is given as:
$\Phi_{ij}^{[l]}(x^{[l-1]}_{j})=w^{[l]}_{ij}(\text{silu}(x^{[l-1]}_{j})+\sum\limits_{m=1}^{G+k-1}c^{[l]}_{ijm}B_{ijm}(x^{[l-1]}_{j})).$
@@
The weights $\{([c^{[l]}_{ijm}],[w^{[l]}_{ij}])\}_{l=1}^{L}$ are learned layerwise, similar to MLPs, by minimizing by final loss for a supervised learning task. However, the bases for the activations could be Fourier (Mehrabian et al., 2024; Xu et al., 2024), wavelet basis functions (Bozorgasl & Chen, 2024), fractals (Yang & Wang, 2025) etc.
@@
Multi-head self-attention (MHSA) in ViTs (Dosovitskiy et al., 2020). The encoder-only vanilla ViT architecture is inspired by the Transformer proposed in (Vaswani et al., 2017). For simplicity, we do not mention the details of the layer normalization and other technicalities. Our central focus is the MHSA architecture.
@@
Let ${X\in\mathbb{R}^{H\times W\times C}}$ be an input image of resolution ${H\times W}$. ViT uses an $p\times p$ patch and generates $N=\tfrac{HW}{p^{2}}$ input tokens. Each token is vectorized along $C$ channels to produce a matrix ${X_{T}\in\mathbb{R}^{N\times p^{2}C}}$, where each row, $X_{T_{i,:}}$, represents a flattened patch across $C$ channels. A learnable embedding matrix, $E\in\mathbb{R}^{p^{2}C\times d}$ projects each row ${X_{T}}_{i,:}$ to an embedding vector in $\mathbb{R}^{d}$ such that After appending the ground-truth class, $x_{{\cal G}}$ and adding the positional encoding matrix, $P_{E}\in\mathbb{R}^{(N+1)\times d}$ of compatible size, the final input matrix, $X_{I}=[x_{{\cal G}};X_{E}]+P_{E}$. The matrix $X_{I}$ is layer normalized and further projected on the row spaces of three learnable weight matrices, $W_{Q},W_{V},W_{k}\in\mathbb{R}^{d\times d}$ to generate the query, key, and value matrices, $Q,K,$ and $V$, respectively, via $Q=X_{I}W_{Q},K=X_{I}W_{K},\;{\rm and}\;V=X_{I}W_{V}.$ These matrices are further divided into $h$ partitions (also, called heads), ${Q^{i},K^{i},V^{i}\in\mathbb{R}^{N\times\tfrac{d}{h}}}$, such that each partition generates a self-attention matrix for the $i^{\rm th}$ head, ${{\cal A}^{i}=\tfrac{Q^{i}{K^{i}}^{\top}}{\sqrt{d/h}}.}$ Finally, we project $V^{i}$ onto the column space of $\sigma({\cal A}^{i})$ as $\sigma({\cal A}^{i})V^{i}$, where $\sigma$ is the Softmax operator applied row-wise. The outputs of different heads are further concatenated into a large matrix ${\begin{bmatrix}\sigma({\cal A}^{1})V^{1}&\sigma({\cal A}^{2})V^{2}&\cdots&\sigma({\cal A}^{h})V^{h}\end{bmatrix}W^{O},}$ post-multiplied by the output weight matrix $W^{O}\in\mathbb{R}^{d\times d}$. This process is sequentially performed in $L$ encoder blocks; see Figure 2(a).

### v2

@@ Abstract
Kolmogorov-Arnold networks (KANs) are a remarkable innovation consisting of learnable activation functions with the potential to capture more complex relationships from data. Presently, KANs are deployed by replacing multilayer perceptrons (MLPs) in deep networks, including advanced architectures such as vision Transformers (ViTs). Given the success of replacing MLP with KAN, this work designs and asks whether a similar replacement in the attention can bring benefits. In this paper, we design the first learnable attention called Kolmogorov-Arnold Attention (KArAt) for ViTs that can operate on any basis, ranging from Fourier, Wavelets, Splines, to Rational Functions. However, learnable activations in the attention cause a memory explosion. To remedy this, we propose a modular version of KArAt that uses a low-rank approximation. By adopting the Fourier basis into this, Fourier-KArAt and its variants, in some cases, outperform their traditional softmax counterparts, or show comparable performance on CIFAR-10, CIFAR-100, and ImageNet-1K datasets. We also deploy Fourier KArAt to ConViT and Swin-Transformer, and use it in detection and segmentation with ViT-Det. We dissect these architectures’ performance on the classification task by analyzing their loss landscapes, weight distributions, optimizer path, attention visualization, and transferability to other datasets, and contrast them with vanilla ViTs. KArAt’s learnable activation shows a better attention score across all ViTs, indicating better token-to-token interactions, contributing to better inference. Still, its generalizability does not scale with larger ViTs. However, many factors, including the present computing interface, affect the relative performance of parameter- and memory-heavy KArAts. We note that the goal of this paper is not to produce efficient attention or challenge the traditional activations; by designing KArAt, we are the first to show that attention can be learned and encourage researchers to explore KArAt in conjunction with more advanced architectures that require a careful understanding of learnable activations. Our open-source code and implementation details are available on: subhajitmaity.me/KArAt
@@ 1 Introduction
Figure 1: (a) Model parameters vs. Top-1 Test accuracy in ImegeNet-1K training of vanilla ViTs [24], Vision KAN (DeiT+KAN) by [17], ViT+KAN and Kolmogorov-Arnold Transformer (KAT) by [80].(b-i) The traditional softmax attention. (b-ii) The Kolmogorov-Arnold Attention (KArAt) replaces the softmax with a learnable operator, $\Phi^{i,j}$. (b-iii) Regular KArAt uses an operator matrix, $\Phi^{i,j}$ with $N^{2}$ learnable units acting on each row of ${\cal A}^{i,j}$, and is prohibitively expensive. (b-iv) Modular KArAt uses an operator $\widehat{\Phi}^{i,j}\in\mathbb{R}^{N\times r}$ with $r\ll N$, followed by a learnable linear projector $W\in\mathbb{R}^{r\times N}$.
@@
Artificial general intelligence has become a rapidly growing research direction, and Kolmogorov-Arnold Network (KAN) [50] marks a remarkable innovation in that. KANs with learnable activation functions can potentially capture more complex relationships and facilitate meaningful interaction between the model and human intuition. After training a KAN on a specific problem, researchers can extract the learned univariate functions that the model uses to approximate complex multivariable functions. By studying these learned functions, researchers can gain insights into the underlying relationships from the data and refine the model. KANs exhibit state-of-the-art performance in finding symbolic function representations [81], continual learning of one-dimensional functions [51].
@@
KANs were integrated with neural network architectures, the primary of them being conventional MLPs or convolution neural networks (CNNs). E.g., [28, 9, 25, 1, 75] combine KAN with CNNs, [44] combine KAN with U-Net, etc. Interestingly, [81] claimed to make the first fairer comparison between KANs and MLPs on multiple ML tasks on small-scale datasets. By setting the same parameter counts in both networks, [81] concluded that KANs with B-Spline basis significantly underperform compared to their MLP counterparts in all ML tasks except symbolic formula representation; [5] reinstates a similar claim for vision. However, KAN’s exploration of more advanced architectures, such as Transformers, remains limited. [17] replace the MLP layers inside the encoder blocks of data-efficient image Transformers (DeiTs) [73] with KANs and proposed DeiT+KAN, [80] proposed two variants: ViT+KAN that replaces the MLP layers inside ViT’s encoder blocks, and Kolmogorov-Arnold Transformer (KAT), albeit similar to ViT+KAN but with a refined group-KAN strategy.
@@
While KANs were claimed to be not superior to simple MLPs in vision tasks [81, 5], we were curious to find out how ViTs can see images through the lens of KANs. In that effort, Figure 1(a) demonstrates the performance of vanilla ViTs (-Tiny, -Small [76, 68], and -Base [24]) and their KAN counterparts on ImageNet-1K [21] training. While different variants of DeiT+KAN [17] and ViT+KAN [80] show a 5%–18% drop in the Top-1 accuracy compared to vanilla ViTs and DeiTs in ImageNet-1K, the KATs [80] show approximately 1%–2.5% gain in Top-1 accuracy compared to vanilla ViTs and DeiTs by keeping about the same number of parameters. These results show KANs require further investigation in more advanced architectures such as Transformers. Nevertheless, first, we want to understand why there is a discrepancy in these two seemingly equivalent implementations.
@@
In search of an answer, we realized that while DeiT+KAN and ViT+KAN replace the MLP layer with KAN in the Transformer’s encoder block, KAT implements a sophisticated group-KAN strategy that reuses a learnable function among a group of units in the same layer and chooses different bases for different encoder blocks. This strategy helps improve the accuracy by keeping almost similar parameter counts. Therefore, simply replacing MLPs with KANs might not guarantee better performance, but a properly designed KAN could. This perspective on the KANs’ performance in conjunction with Transformers opens many research questions. Importantly, we also note that the community remains put from designing a learnable multihead attention module, the heart of the Transformers. Therefore, in the rise of second-generation Transformers, such as Google’s TITAN [6] and SAKANA AI’s Transformer2 [70] that mimic the human brain, we want to ask: Is it worth deploying learnable multi-head self-attention to the (vision) Transformers?
@@
Finally, we note that there is a line of research that proposes efficient attention mechanisms in terms of sparse attention [63, 82, 67, 39, 85, 83, 31] or linear/kernelized attention [32, 58, 57, 56, 52, 55], to remedy computational and memory complexities of attention calculation. In contrast, this paper tries to understand how learnable multi-head self-attention (MHSA) modules can perform over regular self-attention used in ViTs, a defining technology in vision in the last six years. Our contributions are:
@@
Designing a Learnable Attention Module for ViTs. Since language is significantly more complicated, less structured, and has a higher information entropy than an image, we limit this pilot study to the ViTs [24] for image classification. In §3, we propose a general learnable Kolmogorov-Arnold multi-head self-attention, or KArAt, for ViTs that can operate on any basis. Due to the memory explosion, any learnable MHSA operator matrix in ideal KArAt causes a training bottleneck. We propose a low-rank approximation of the operator matrix to reduce memory requirements to a feasible level, but leave the possibility of exploring alternative techniques. By adopting the Fourier basis into this modular version, we design Fourier KArAt and its variants; see §4 for details.
@@
Benchmarking, Evaluation, and Analysis (§56). We benchmark Fourier KArAt on CIFAR-10, CIFAR-100, [40] and ImageNet-1K [21] datasets and evaluate their performance against vanilla ViTs (ViT-Tiny, -Small [76, 68], and -Base [24]), and other popular ViTs (ConViT [20] and Swin-Transformer [49]). We use small-scale datasets, SVHN [54], Oxford Flowers 102 [59], and STL-10 [18] for understanding the transfer learning capability of the models. Additionally, we embed KArAt with ViT-Det [46] for the object detection and segmentation task on MS COCO [48]. We dissect KArAts’ performance and generalization capacity by analyzing their loss landscapes, weight distributions, optimizer path, and attention visualization, and compare them with traditional softmax attention in vanilla ViTs.
@@ 2 Background
Multi-layer Perceptrons (MLPs) consist of $L$ layers. By convention, $a^{[0]}=x^{[0]}$ denotes the input data and the $l^{\rm th}$-layer is defined for input $x^{[l-1]}\in\mathbb{R}^{d_{l-1}}$ as
@@
In MLP, the nonlinear activation functions, $\Phi(\cdot)$, are fixed [7]. For supervised tasks, given a training dataset $D$ with $N$ elements of the form (input, ground-truth) pairs, $\{(x_{i},y^{\star}_{i})\}_{i=1}^{N}$, the loss, $\mathcal{L}(X,Y^{\star}|\mathcal{W})=d_{Y}(a^{[L]},Y^{\star})$ (metric induced by the space $\mathbb{R}^{d_{L}}$) is calculated in the forward pass. The layerwise weights, $\{(W_{l},b_{l})\}_{l\in[L]}$ are learned by minimizing $\mathcal{L}(X,Y^{\star}|\mathcal{W})$.
Kolmogorov-Arnold Network (KAN) [51] is a neural network involving learnable activations parametrized by a chosen set of basis functions defined on a set of grid points or knots. The idea for this network stems from the Kolmogorov-Arnold Representation Theorem.
@@
[38] For any multivariate continuous function, $f:[0,1]^{n}\to\mathbb{R}$, there exists a finite composition of continuous single-variable functions, $\phi_{q,p}:[0,1]\to\mathbb{R},\Phi_{q}:\mathbb{R}\to\mathbb{R}$ such that
$f(x)=f(x_{1},x_{2},\cdots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right).$
@@
Theorem 1 describes an exact finite expression using 2 layers, but [51] generalized this representation to multi-layers via KAN; it is analogous to the Universal Approximation Theorem [34]. For input $x^{[0]}$, an $L$-layer KAN is given as $\text{KAN}(x^{[0]})=\Phi^{L}\circ\cdots\circ\Phi^{l}\circ\cdots\circ\Phi^{1}(x^{[0]})$. In each KAN layer, the activation function $\Phi$ is learnable, and ${\Phi^{[l]}=[\Phi_{ij}^{[l]}]}$ operates on ${x^{[l-1]}\in\mathbb{R}^{d_{l-1}}}$ to produce ${x^{[l]}\in\mathbb{R}^{d_{l}}},$ such that:
${x^{[l]}_{i}=\Phi^{[l]}_{i:}(x^{[l-1]})=\sum_{j=1}^{d_{l-1}}\Phi^{[l]}_{ij}(x^{[l-1]}_{j})}.$
@@
Originally, [51] chose the B-spline basis functions. That is, the activation functions were defined to be $\phi(x)=w(b(x)+\text{spline}(x))$, where $b(x)=\texttt{SiLU}(x)=\frac{x}{1+e^{-x}}$ and $\text{spline}(x)=\sum\limits_{i}c_{i}B_{i}(x)$, with $B_{i}(\cdot)$ being one of the $k$-th degree B-spline basis functions parametrized by the $G$ grid points on a uniform grid, $[-I,I]$. In this case, the representation for each function is given as:
$\Phi_{ij}^{[l]}(x^{[l-1]}_{j})=w^{[l]}_{ij}(\texttt{SiLU}(x^{[l-1]}_{j})+\sum\limits_{m=1}^{G+k-1}c^{[l]}_{ijm}B_{ijm}(x^{[l-1]}_{j})).$ (1)
@@
The weights ${\{([c^{[l]}_{ijm}],[w^{[l]}_{ij}])\}_{l=1}^{L}}$ are learned by minimizing the loss for a supervised learning task. The bases for KANs’ activations could be Fourier [53, 78], wavelet [10], fractals [80], etc.
@@
Multi-head self-attention (MHSA) in ViTs [24]. The encoder-only vanilla ViT architecture is inspired by the Transformer proposed in [74]. For simplicity, we do not mention the details of the layer normalization and other technicalities. Our central focus is the MHSA architecture.
@@
Let ${X\in\mathbb{R}^{H\times W\times C}}$ be an input image of resolution ${H\times W}$. ViT uses a $p\times p$ patch and generates $N=\tfrac{HW}{p^{2}}$ input tokens. Each token is vectorized along $C$ channels to produce ${X_{T}\in\mathbb{R}^{N\times p^{2}C}}$, where each row, ${X_{T}}_{i,:}$, represents a flattened patch across channels. A learnable embedding matrix, $E\in\mathbb{R}^{p^{2}C\times d}$ projects ${X_{T}}_{i,:}$ to an embedding vector in $\mathbb{R}^{d}$ such that after appending the learnable class token, $x_{{\cal G}}$ and adding the positional encoding matrix, $P_{E}\in\mathbb{R}^{(N+1)\times d}$ of compatible size, the final input, $X_{I}=[x_{{\cal G}};X_{E}]+P_{E}$. The matrix $X_{I}$ is layer normalized and further projected on the row spaces of three learnable weight matrices, $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$ to generate the query, key, and value matrices, $Q,K,$ and $V$, respectively, via $Q=X_{I}W_{Q},K=X_{I}W_{K},\;{\rm and}\;V=X_{I}W_{V}$. These matrices are further divided into $h$ partitions (also, called heads), ${Q^{i},K^{i},V^{i}\in\mathbb{R}^{N\times\tfrac{d}{h}}}$, such that each partition generates a self-attention matrix for the $i^{\rm th}$ head, ${{\cal A}^{i}=\tfrac{Q^{i}{K^{i}}^{\top}}{\sqrt{d/h}}.}$ Finally, we project $V^{i}$ onto the column space of $\sigma({\cal A}^{i})$ as $\sigma({\cal A}^{i})V^{i}$, where $\sigma$ is the softmax operator applied row-wise. The outputs of different heads are further concatenated into a large matrix ${\begin{bmatrix}\sigma({\cal A}^{1})V^{1}&\sigma({\cal A}^{2})V^{2}&\cdots&\sigma({\cal A}^{h})V^{h}\end{bmatrix}W^{O},}$ post-multiplied by the learnable weight, $W^{O}\in\mathbb{R}^{d\times d}$. This process is sequentially performed in $L$ encoder blocks; see Figure 1(b)(i).